Problem 3 Engineers often bank highway curves to reduce the vehicle’s reliance on friction while navigating turns. If the bank is angled properly, the posted speed limit is effective in both wet and dry weather. For a curve of radius 13 0m and a speed limit of 100 km/hr, what is the minimum bank angle such that a reliance on friction is not necessary?
Answer #1

in vertical, N cos(theta) - m g = 0

N cos(theta) = m g ....(i)

in horizontal, N sin(theta) = m v^2 / r .... (ii)

(ii)/(i) => tan(theta) = v^2 / r g

v= 100 km/h = 100 x 1000 / 3600 m/s = 27.8 m/s

theta = tan^-1(27.8^2 / (13 x9.8))

= 80 deg